# Dialetto, ma Quanto Dialetto? Transcribing and Evaluating Dialects on a Continuum

### v1

@@ Abstract
There is increasing interest in looking at dialects in NLP. However, most work to date still treats dialects as discrete categories. For instance, evaluative work in variation-oriented NLP for English often works with Indian English or African-American Venacular English as homogeneous categories (Faisal et al., 2024; Ziems et al., 2023), yet even within one variety there is substantial variation. We examine within-dialect variation and show that performance critically varies within categories. We measure speech-to-text performance on Italian dialects, and empirically observe a geographical performance disparity. This disparity correlates substantially (-0.5) with linguistic similarity to the highest performing dialect variety. We cross-examine our results against dialectometry methods, and interpret the performance disparity to be due to a bias towards dialects that are more similar to the standard variety in the speech-to-text model examined. We additionally leverage geostatistical methods to predict zero-shot performance at unseen sites, and find the incorporation of geographical information to substantially improve prediction performance, indicating there to be geographical structure in the performance distribution.
[…]
booktabs
@@ 1 Introduction
An increasing body of work in Natural Language Processing (NLP) has called attention to the disparity in research focus between high-resource, standardized linguistic varieties and empirical linguistic variation (Plank, 2016; Hupkes et al., 2023; Kantharuban et al., 2023; Chang et al., 2024). While there are many types of variation (e.g. genre, register), dialect variation has emerged as a particular point of focus, with increasing availability of evaluative benchmarks (Faisal et al., 2024; Ziems et al., 2023), dialect-specific datasets (Dogan-Schönberger et al., 2021; Blaschke et al., 2024), and methodological contributions (Blaschke et al., 2023; Demszky et al., 2021) towards dialect-robust models (Zampieri et al., 2020).
@@
A considerable amount of such work conceptualizes dialects solely as discrete linguistic categories that stand side-by-side with the standard variety (e.g. African-American Vernacular English vs. mainstream American English) (Faisal et al., 2024; Ziems et al., 2023). However, prior work in dialectology has noted that dialect relations often stand in a continuum, where similarity between varieties slowly decreases the further away from a given geographical site, rather than being a sharp transition (Heeringa and Nerbonne, 2001). For dialect NLP, this means that a purely discrete conceptualization of linguistic categories not only largely overlooks the dialect continuum, but also leaves important evaluation gaps. As consequence, gradient variation within the category may not be evenly described (Jones, 2015; Labov, 2012), and lesser known transitional varieties (Jeszenszky et al., 2018) between the linguistic categories examined may be left out of evaluative benchmarks.
@@
Figure 1: Procedure through which we obtain our speech-to-text scores. Semantically-equivalent sentence audios are sampled across geographically-balanced dialects, which are then transcribed to the standard variety with ASR models, and evaluated with machine translation evaluation metrics.
@@
We perform a more fine-grained investigation of such a regional gap in speech-to-text models to gain insights on cross-lingual transfer performance by drawing upon geostatistical and dialectometric techniques. In contrast with prior work which evaluates performance disparity on a discrete basis (e.g. Chilean Spanish, Argentinian Spanish) (Kantharuban et al., 2023), we place dialects on a continuum using a large-scale geotagged dataset of related Italian dialects, where we conceptualize dialect relations to stand in a continuum (Heeringa and Nerbonne, 2001). As such, we perform zero-shot speech-to-text on semantically-equivalent speech samples across geographically-contiguous dialect sites, where in line with Kantharuban et al. (2023), we find evidence of zero-shot performance of speech-to-text models on dialects correlating with similarity to the standard variety. Our contributions are as follows:
@@ 2.1 Italian Dialect Dataset
We conduct our study on Vivaldi (Tosques and Castellarin, 2013), a geo-tagged parallel corpus of spoken Italian dialect varieties in audio form. The corpus contains data at different levels of linguistic units (e.g. word, sentence, discourse). The data is collected across 293 sites in Italy, and is divided into the categories of phonetic, lexical, morphological, syntactic, and discourse level data. Our criterion of using semantically-equivalent data across dialect sites is motivated by the fact that there are syntactic, lexical, and phonetic differences between dialects, which serves as a more realistic basis for evaluating zero-shot performance on dialects. As such, we leverage 15 sentence-level recordings and 20 word-level recordings per site in our experiments. To ensure a fair comparison between sites and because of the rich linguistic variety observed in Italy Ramponi (2024), we further filter for only dialect sites that fall under the Italic language branch according to the metadata, thereby removing data from Bavarian, Greek, Occitan, among others. This results in 223 sites.
In view of the scale of variation in Italian dialects, where differences between groups may warrant the status of individual languages, we also evaluate our results on only the Tuscan subgroup, which qualitative work establishes as bearing the most similarity to standard Italian. Table 1 summarizes the statistical information of our data.
@@
[…]
Table 1: Statistics on the subset of Vivaldi (Tosques and Castellarin, 2013), the Italian dialect dataset used in our experiments.
[…]
@@ 3.1 Speech-to-Text
For our speech-to-text model, we employ Whisper (Radford et al., 2023), a family of encoder-decoder speech-to-text models. Our experiment utilizes Whisper-large-v3, which is trained on 1 million hours of weakly-labeled and 4 million hours of pseudo-labeled audio data. The training regime for Whisper is both multilingual and multi-task, where samples are either asked to be transcribed into the original language, or to be translated into English. This is achieved by way of special tokens (e.g. <lang>, <translate>, <transcribe>). In addition, the data format for long-form transcriptions includes a token <prev> to denote the previous context during training. At inference time, the space of this prior context can be used to achieve prompting, where the ASR output would be conditioned on this prior context. We take advantage of this prior context to prompt the model to transcribe the speech in standard language (e.g. "Questa è una frase italiana: "; English translation: "this is an Italian sentence: "). This is necessary due to the spoken nature of dialect varieties in Italy, where there is often no widely used written variety that corresponds with what is spoken, and speakers of such varieties would write often only in the standard variety.
@@ 3.2 Evaluation
Dolev et al. (2024) report Whisper as a viable system for speech-based dialect-to-standard ASR for Swiss German when transcribed to standard German text and evaluated with BLEU (Papineni et al., 2002). Given a similar mismatch between input and output for dialect speech to standard text for our Italian dialect data, we follow Dolev et al. (2024) in evaluating our dialect speech-to-text output with standard machine translation evaluation metrics. We employ BLEU (Papineni et al., 2002) and chrF (Popović, 2015) [^1] We employ the implementation in SacreBLEU (Post, 2018)., both of which are based on ngram overlap. Due to the expected mismatch between Italian dialect speech and standard Italian text, we increase the number of gold references to allow for more opportunity for alternative yet valid phrasings to be counted as correct. We follow prior work in expanding the number of gold references by way of a LLM-based paraphrasing approach (Tang et al., 2024; Zeng et al., 2024). Table 2 gives an example of the original and generated gold references for standard Italian. In our experiments, we generate 10 additional references per item [^2] We use Meta-Llama-3.1-70B-Instruct: https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, in addition to the original gold standard. Note that we do not employ standard ASR metrics such as word error rate or character error rate, as we do not expect the speech and the text to align well for every variety, due to the non-written nature of non-standard varieties.
3.3 Dialectometric Analysis
To quantify the linguistic distance between varieties, we follow established approaches in dialectometry (Wieling and Nerbonne, 2015) for both distance computation and visualization, which we employ on 20 audio recordings of words per site.
3.3.1 Linguistic Distance
To quantify the linguistic distance between dialect varieties, we follow Bartelds and Wieling (2022) in adopting self-supervised speech representations for the extraction of features on semantically equivalent words, upon which dynamic time warping (DTW) can be applied for a measure of phonetic distance. Formally, let $X$ and $Y$ denote two lists of semantically equivalent words from two dialect varieties, where $X=[w_{X}^{1},w_{X}^{2},\dots,w_{X}^{n}]$ and $Y=[w_{Y}^{1},w_{Y}^{2},\dots,w_{Y}^{n}]$, with $w_{X}^{i}$ and $w_{Y}^{i}$ representing semantically equivalent items. For each pair of items $(w_{X}^{i},w_{Y}^{i})$, we extract their acoustic features using XLSR-53, a self-supervised speech representation model, where we employed an off-the-shelf finetune on Italian data in the Common Voice (Ardila et al., 2019) dataset [^3] https://huggingface.co/jonatasgrosman/wav2vec2-large-xlsr-53-italian. Let $\mathbf{f}(w_{X}^{i})$ and $\mathbf{f}(w_{Y}^{i})$ represent the feature vectors for the words $w_{X}^{i}$ and $w_{Y}^{i}$, respectively. The acoustic distance $dist(w_{X}^{i},w_{Y}^{i})$ for the $i$-th pair is then computed by way of dynamic time warping, which obtains the distance between two time series that may vary in speed and length by computing the shortest path in a cost matrix. Following Bartelds and Wieling (2022), the distance is length normalized for a fair comparison between sites. The acoustic distance $Dist(X,Y)$ between dialect varieties $X$ and $Y$ is then computed by averaging the pairwise distances across all $n$ items in the lists:
$Dist(X,Y)=\frac{1}{n}\sum_{i=1}^{n}dist(w_{X}^{i},w_{Y}^{i})$
@@ 3.3.1 Linguistic Distance
[…]
3.3.2 Multidimensional Scaling
Nerbonne (2010) shows that a map depicting dialect relations as a continuous surface can be achieved by leveraging (classical) multidimensional scaling [^4] We use the implementation in dialectR (Shim and Nerbonne, 2022).. Multidimensional scaling (MDS) is a dimensionality reduction method, which takes as input a distance matrix and aims to project it to a lower-dimensional space while aiming to preserve the distances in the original high-dimensional space. Formally, given data points $X=\{x_{1},x_{2},\dots,x_{n}\}$, let $D=[d_{ij}]$ be the distance matrix, where $d_{ij}$ represents the distance between $x_{i}$ and $x_{j}$. MDS seeks to minimize the following objective, termed the stress function:
$S(Y)=\sqrt{\frac{\sum_{i,j}(d_{ij}-\|y_{i}-y_{j}\|)^{2}}{\sum_{i,j}d_{ij}^{2}}}$
@@ 3.4 Geostatistical Analysis
Prior work has found geographical proximity between pivot and target language to be an important predictor of cross-lingual transfer (Ahuja et al., 2022; Samardžić et al., 2022; Lin et al., 2019). The varieties examined in Ahuja et al. (2022) cover the span of languages, where the performance is argued to be due to overlap in typological and vocabulary overlap. We propose that for varieties within the same language, the geographical signal is arguably even stronger due to higher geographical proximity and fewer confounding factors (Shim et al., 2024; Jeszenszky et al., 2017), such that the signal can be helpful in predicting zero-shot performance for varieties within a given language. To verify this claim, we leverage geostatistical techniques to predict the zero-shot performance of Whisper on unseen held-out sites. We employ three geostatistical interpolation methods in our experiments [^5] We use implementations in the R package gstat (Pebesma, 2004; Gräler et al., 2016): nearest neighbor interpolation (NN), inverse distance weighting (IDW), and kriging. While NN and IDW do not make assumptions on the geographical distribution of the data, the use of kriging makes the assumption of stationarity, which is the assumption that the mean and variance are constant across space. Given that the variable we model is speech-to-text performance on dialects, its geographical distribution arguably may violate the assumption of stationarity, depending on how well the model generalizes to the dialect varieties in different regions. Given such a potential for non-stationarity in the data, we expect this to require modelling in order to better satisfy the assumption. We detail our treatment of this in Section 3.4.5.
In our experiments, we perform an 80/10/10 split for training, validation, and test data, where we tune hyperparameters by way of a grid search on the validation set, and report the RMSE on the test set. In addition, we measure the impact of training data size across the geostatistical methods examined, where we sample the training data from a percentage range of 0.1 to 1.0. For each point in the percentage range, we repeat the sampling procedure 100 times and take the mean of the interpolation RMSE across the 100 runs, in order to ensure that the performance reported is representative of the data. Table 4 summarizes our results; Figure 3 shows the effect of training data size.
[…]
Table 4: RMSE of geostatistical interpolation methods on unseen sites.
@@ 3.4.1 Baseline
To verify the extent to which increasing levels of geospatial information is helpful for prediction, we employ nearest neighbor interpolation as a baseline, where the predicted value of a sample is taken to be identical to its geographically nearest neighbor in the training data.
@@ 3.4.2 Inverse Distance Weighting
Inverse distance weighting (IDW) is a geostatistical interpolation method, where the estimated value at a target point $\hat{v}(x_{0})$ is computed as a weighted average of the known values from surrounding data points. The weights are inversely proportional to the distance from the target point, with closer points having more influence. Formally, given a set of known points $x_{1},x_{2},\dots,x_{n}$ with corresponding values $v(x_{1}),v(x_{2}),\dots,v(x_{n})$, the interpolated value at $x_{0}$ is defined as:
@@ 3.4.3 Variogram
A variogram is a method in geostatistics that is used to quantify the degree of spatial autocorrelation between data points. In contrast to IDW, which assumes a deterministic decrease in similarity as distance increases, a variogram provides a probabilistic approach to measuring how spatial correlation between values changes with increasing separation distance, allowing for a more data-driven approach towards deriving the weights of the values used to interpolate the value at an unknown site. Formally, the variogram $\gamma(h)$ is defined as:
@@
The function $\gamma(h)$ estimates the spatial variance as a function of distance, with larger values of $h$ indicating weaker correlation between points. A variogram is typically visualized by plotting $\gamma(h)$ against $h$, which allows for a visual interpretation of the y-intercept, representing measurement error or spatial variability at very short distances (the nugget); the point at which spatial correlation diminishes (the sill), and the distance at which the variogram reaches the sill, beyond which points are effectively uncorrelated (the range). A function is typically then fit to the empirical values by adjusting for the parameters of nugget, sill, and range, in order for a continuous model to be obtained that best fits the empirical distribution of the data. Such a model of how the variance varies with distance forms the basis for kriging, a more sophisticated interpolation method that we employ in our experiments and describe next. In our experiments, we automatically fit the theoretical variogram by way of the best least squares fit to the data.

### v2

@@ Abstract
There is increasing interest in looking at dialects in NLP. However, most work to date still treats dialects as discrete categories. For instance, evaluative work in variation-oriented NLP for English often works with Indian English or African-American Venacular English as homogeneous categories, yet even within one variety there is substantial variation. We examine within-dialect variation and show that performance critically varies within categories. We measure speech-to-text performance on Italian dialects, and empirically observe a geographical performance disparity. This disparity correlates substantially (-0.5) with linguistic similarity to the highest performing dialect variety. We cross-examine our results against dialectometry methods, and interpret the performance disparity to be due to a bias towards dialects that are more similar to the standard variety in the speech-to-text model examined. We additionally leverage geostatistical methods to predict zero-shot performance at unseen sites, and find the incorporation of geographical information to substantially improve prediction performance, indicating there to be geographical structure in the performance distribution.
@@ 1 Introduction
An increasing body of work in Natural Language Processing (NLP) has called attention to the disparity in research focus between high-resource, standardized linguistic varieties and empirical linguistic variation (Plank, 2016; Kantharuban et al., 2023; Chang et al., 2024). While there are many types of variation (e.g. genre, register), dialect variation has emerged as a particular point of focus, with increasing availability of evaluative benchmarks (Faisal et al., 2024; Ziems et al., 2023), dialect-specific datasets (Dogan-Schönberger et al., 2021; Blaschke et al., 2024), and methodological contributions (Blaschke et al., 2023; Demszky et al., 2021) towards dialect-robust models (Zampieri et al., 2020).
@@
A considerable amount of work conceptualizes dialects solely as discrete linguistic categories that stand side-by-side with the standard variety (e.g. African-American Vernacular English vs. mainstream American English) (Faisal et al., 2024; Ziems et al., 2023). However, prior work in dialectology has noted that dialect relations often stand in a continuum, where similarity between varieties slowly decreases the further away they are from a given geographical site, rather than being a sharp transition (Heeringa and Nerbonne, 2001). For dialect NLP, this means that a purely discrete conceptualization of linguistic categories not only largely overlooks the dialect continuum, but also leaves important evaluation gaps, which can even lead to social harm. This is because gradient variation within the category may not be evenly described (Jones, 2015; Labov, 2012), and lesser known transitional varieties (Jeszenszky et al., 2018) between the linguistic categories examined may be left out of evaluative benchmarks.
@@
Figure 1: Procedure through which we obtain our speech-to-text scores. Semantically-equivalent sentence audios are sampled across geographically-balanced dialects, which are then transcribed to the standard variety with speech-to-text models, and evaluated with machine translation evaluation metrics.
Concretely, this paper addresses two research questions:
  1. RQ1: Is the distribution of speech-to-text performance on dialect speech a geographically autocorrelated variable?
  2. RQ2: To what extent is the distribution predictable by way of phonetic similarity to the best performing variety?
Our first question (RQ1) is motivated by prior work in the geosciences, which has consistently relied on the insight that “near things are more related than distant things”—a principle known as Tobler’s first law of geography (Tobler, 1970)—for interpolation of missing values (Matheron, 1963). On the other hand, our second question (RQ2) is motivated by the possibility of pretraining data containing differing amounts of dialect speech, which is likely to serve as a confounding variable to the correlation between phonetic similarity and speech-to-text performance.
@@
To answer the two questions, we perform a fine-grained investigation of such a regional gap in speech-to-text models to gain insights on cross-lingual transfer performance by drawing upon geostatistical and dialectometric techniques. In contrast with prior work which evaluates performance disparity on a discrete basis (e.g. Chilean Spanish, Argentinian Spanish) (Kantharuban et al., 2023), we place dialects on a continuum using a large-scale geotagged dataset of related Italian dialects, where we conceptualize dialect relations to stand in a continuum (Heeringa and Nerbonne, 2001). As such, we perform zero-shot speech-to-text on semantically-equivalent speech samples across geographically-contiguous dialect sites, where in line with Kantharuban et al. (2023), we find evidence of zero-shot performance of speech-to-text models on dialects correlating with similarity to the standard variety. [^1] Our code is available here: https://github.com/mainlp/dialetto Our contributions are as follows:
@@ 2.1 Italian Dialect Dataset
We conduct our study on Vivaldi (Tosques and Castellarin, 2013), a geo-tagged parallel corpus of spoken Italian dialect varieties in audio form. The corpus contains data at different levels of linguistic units (e.g. word, sentence, discourse). The data is collected across 293 sites in Italy, and is divided into the categories of phonetic, lexical, morphological, syntactic, and discourse level data. Our criterion for using semantically-equivalent data across dialect sites is motivated by the fact that there are syntactic, lexical, and phonetic differences between dialects, which serves as a more realistic basis for evaluating zero-shot performance on dialects. As such, we leverage 15 sentence-level recordings and 20 word-level recordings per site in our experiments. To ensure a fair comparison between sites and because of the rich linguistic variety observed in Italy Ramponi (2024), we further filter for only dialect sites that fall under the Italic language branch according to the metadata, thereby removing data from Bavarian, Greek, Occitan, among others. This results in 223 sites. Table 1 summarizes the statistical information of our data.
In view of the scale of variation in Italian dialects, where differences between groups may warrant the status of individual languages, we also evaluate our results on only the Tuscan subgroup, which qualitative work establishes as bearing the most similarity to standard Italian (Wieling et al., 2014).
@@
rowsep=0pt # dialects used # words used # sentences used 223 20 15
Table 1: Statistics on the subset used of Vivaldi (Tosques and Castellarin, 2013), an Italian dialect corpus.
@@ 3.1 Speech-to-Text
For our speech-to-text model, we employ Whisper (Radford et al., 2023), a family of encoder-decoder speech-to-text models. Our experiment utilizes Whisper-large-v3, which is trained on 1 million hours of weakly-labeled and 4 million hours of pseudo-labeled audio data. The training regime for Whisper is both multilingual and multi-task, where samples are either asked to be transcribed into the original language, or to be translated into English. This is achieved by way of special tokens (e.g. <lang>, <translate>, <transcribe>). In addition, the data format for long-form transcriptions includes a token <prev> to denote the previous context during training. At inference time, the space of this prior context can be used to achieve prompting, where the speech-to-text output would be conditioned on this prior context. We take advantage of this prior context to prompt the model to transcribe the speech in standard language (e.g. "Questa è una frase italiana: "; English translation: "this is an Italian sentence: "). This is necessary due to the spoken nature of dialect varieties in Italy, where there is often no widely used written variety that corresponds with what is spoken, and speakers of such varieties would write often only in the standard variety.
@@ 3.2 Evaluation
Dolev et al. (2024) report Whisper as a viable system for speech-based dialect-to-standard speech-to-text for Swiss German when transcribed to standard German text and evaluated with BLEU (Papineni et al., 2002). Given a similar mismatch between input and output for dialect speech to standard text for our Italian dialect data, we follow Dolev et al. (2024) in evaluating our dialect speech-to-text output with standard machine translation evaluation metrics. We employ BLEU (Papineni et al., 2002) and chrF (Popović, 2015) [^2] We employ SacreBLEU (Post, 2018). both of which are based on n-gram overlap. Due to the expected mismatch between Italian dialect speech and standard Italian text, we increase the number of gold references to allow for more opportunity for alternative yet valid phrasings to be counted as correct. We follow prior work in expanding the number of gold references by way of a LLM-based paraphrasing approach (Tang et al., 2024; Zeng et al., 2024). Table 2 gives an example of the original and generated gold references for standard Italian. In our experiments, we generate 10 additional references per item, [^3] We use Meta-Llama-3.1-70B-Instruct: https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct in addition to the original gold standard. Note that we do not employ standard ASR metrics such as word error rate or character error rate, as we do not expect the speech and the text to align well for every variety, due to the non-written nature of non-standard varieties.
@@
rowsep=0pt Original Generated (1) Si munge due volte al giorno Si munge due volte ogni giorno EN: One milks two times per day One milks two times every day (2) Domani tornerò a casa Domani ritornerò a casa mia EN: I will go home tomorrow I will return to my home tomorrow
@@ 3.3 Geostatistical Analysis
In this section, we introduce geostatistical (Matheron, 1963; Cressie, 1989) methods for the interpolation of speech-to-text performance at unseen sites, where geostatistics refers to a family of statistical techniques designed to model spatial data. Our introduction of such methods is driven by two considerations: to understand whether the geographical distribution of model performance is indeed sufficiently autocorrelated for such interpolation to work, and for the practical concern that data collected for dialects may be more sparsely distributed than desired due to difficulties in collection, thus raising the need for interpolation.
Prior work has found geographical proximity between pivot and target language to be an important predictor of cross-lingual transfer (Ahuja et al., 2022; Samardžić et al., 2022; Lin et al., 2019). The varieties examined in Ahuja et al. (2022) cover the span of languages, where the performance is argued to be due to overlap in typological and vocabulary overlap. We propose that for varieties within the same language, the geographical signal is arguably even stronger due to higher geographical proximity and fewer confounding factors (Shim et al., 2024; Jeszenszky et al., 2017), such that the signal can be helpful in predicting zero-shot performance for varieties within a given language. To verify this claim, we leverage geostatistical techniques to predict the zero-shot performance of Whisper on unseen held-out sites. We employ three geostatistical interpolation methods in our experiments: [^4] We use implementations in the R package gstat (Pebesma, 2004; Gräler et al., 2016). nearest neighbor interpolation (NN) (Sibson, 1981), inverse distance weighting (IDW) (Shepard, 1968), and kriging (Oliver and Webster, 1990). While NN and IDW do not make assumptions on the geographical distribution of the data, the use of kriging makes the assumption of stationarity, which is the assumption that the mean and variance are constant across space. Given that the variable we model is speech-to-text performance on dialects, its geographical distribution arguably may violate the assumption of stationarity, depending on how well the model generalizes to the dialect varieties in different regions. Given such a potential for non-stationarity in the data, we expect this to require modelling in order to better satisfy the assumption. We detail our treatment of this in Section 3.3.5.
In our experiments, we perform an 80/10/10 split for training, validation, and test data, where we tune hyperparameters by way of a grid search on the validation set, and report the root mean square error (RMSE) on the test set. In addition, we measure the impact of training data size across the geostatistical methods examined, where we sample the training data from a percentage range of 0.1 to 1.0. For each point in the percentage range, we repeat the sampling procedure 100 times and take the mean of the interpolation RMSE across the 100 runs, in order to ensure that the performance reported is representative of the data. Table 6 summarizes our results; Figure 3 shows the effect of training data size.
@@ 3.3.1 Baseline
To verify the extent to which increasing levels of geospatial information is helpful for prediction, we employ nearest neighbor interpolation (Sibson, 1981) as a baseline, where the predicted value of a sample is taken to be identical to its geographically nearest neighbor in the training data.
@@ 3.3.2 Inverse Distance Weighting
Inverse distance weighting (IDW) (Shepard, 1968) is a geostatistical interpolation method, where the estimated value at a target point $\hat{v}(x_{0})$ is computed as a weighted average of the known values from surrounding data points. The weights are inversely proportional to the distance from the target point, with closer points having more influence. Formally, given a set of known points $x_{1},x_{2},\dots,x_{n}$ with corresponding values $v(x_{1}),v(x_{2}),\dots,v(x_{n})$, the interpolated value at $x_{0}$ is defined as:
@@ 3.3.3 Variogram
Our method of kriging (Oliver and Webster, 1990) relies on the concept of a variogram, where a variogram (Cressie, 1985) is a method in geostatistics that is used to quantify the degree of spatial autocorrelation between data points. In contrast to IDW, which assumes a deterministic decrease in similarity as distance increases, a variogram provides a probabilistic approach to measuring how spatial correlation between values changes with increasing separation distance, allowing for a more data-driven approach towards deriving the weights of the values used to interpolate the value at an unknown site. Formally, the variogram $\gamma(h)$ is defined as:
@@
The function $\gamma(h)$ estimates the spatial variance as a function of distance, with larger values of $h$ indicating weaker correlation between points. A variogram is typically visualized by plotting $\gamma(h)$ against $h$, which allows for a visual interpretation of the y-intercept, representing measurement error or spatial variability at very short distances (the nugget); the point at which spatial correlation diminishes (the sill), and the distance at which the variogram reaches the sill, beyond which points are effectively uncorrelated (the range). A function is then typically fit to the empirical values by adjusting for the parameters of nugget, sill, and range, in order for a continuous model to be obtained that best fits the empirical distribution of the data. Such a model of how the variance varies with distance forms the basis for kriging, a more sophisticated interpolation method that we employ in our experiments and describe next. In our experiments, we automatically fit the variogram by way of the best least squares fit to the data.